culture is grown using acetate as the sole source of carbon and envergy. The cell yield ) carbon will be found as to be 045 (ie. (acetate 045g cell mass was produced per 1g of acetate utilized). What percentage of the ass as CO? You may assume that you begin with
Answer #1

cell yield is estimated to be 0.45 , which means 1g of acetate is utilized for 0.45 g of cell mass.

1 mole(59 g) of acetate is utilized for = 0.45 *59

= 26.55 g of cell mass

= 26.55/113 moles of cell mass

= 0.234 moles of cell mass

1 mole of acetate contain 24 g of carbon which is utilized to produce 0.234 g of cell mass .

1 mole of cell mass contain 60 g of carbon.

0.234 g of cell mass contain = 0.234*60

= 14.09 g of carbon

percentage of acetate carbon found as cell mass = (14.09/24) *100

= 58.74%

Answer is 58.74%